OpenStudy (anonymous):

The variable Z is inversely proportional to X. When X is 3, Z has the value 2. What is the value of Z when X = 13

OpenStudy (anonymous):

@Nnesha can you help?

Nnesha (nnesha):

do you know the definition of inverse proportion ?

Nnesha (nnesha):

i mean formula*

OpenStudy (anonymous):

i know to do it , its just the decimals get me confused. 2/3 is what? then whatver that is i just multiply it by 13

OpenStudy (insa):

As z is inversely proportional to x. Then z=k/x substitute the values of z and k and find the value of k

Nnesha (nnesha):

you don't have to divide Z = k/ z substitute z and k values and x is dividing by k so you need to do opposite of division

OpenStudy (insa):

After finding the value of k, let z=k/x. Then substitute the value of k u found and the other value of k, solve it further and you'll get the answer :)

OpenStudy (anonymous):

so is the answer 78?

Nnesha (nnesha):

k equal to what ?

Nnesha (nnesha):

k is constant of variation first you need to solve for k \[\huge\rm \color{reD}{z}=\frac{ k }{ \color{blue}{x} }\] \[\color{ReD}{z =2}~~~~\color{blue}{x=3}\]

OpenStudy (anonymous):

so then itll be 2=26?

OpenStudy (anonymous):

or is it 6?

Nnesha (nnesha):

yes k=6 now use the same formula substitute k and z values solve for z

OpenStudy (anonymous):

so then its 2=2

Nnesha (nnesha):

how ?

OpenStudy (anonymous):

2=6/3 ?

Nnesha (nnesha):

no you need to find z when x =13 this is 2nd part so substitute 13 for x

OpenStudy (anonymous):

so 2=6/13?

Nnesha (nnesha):

no u need to find z so it should be z =6/13

OpenStudy (anonymous):

0.461?

Nnesha (nnesha):

yeah or leave it as fraction form

Nnesha (nnesha):

.4615 ^round it

OpenStudy (anonymous):

0.462

Nnesha (nnesha):

lO_Oks right

OpenStudy (anonymous):

i need one more The variable Z is directly proportional to X, and inversely proportional to Y. When X is 6 and Y is 3, Z has the value 6. What is the value of Z when X = 10, and Y = 8

Nnesha (nnesha):

directly proportional = y =kx multiply k with x inverse = y = k/x divide k by x so translate that into the equation

Nnesha (nnesha):

is means equal sign so z =

OpenStudy (anonymous):

so multiply 6x6 and 6x3

Nnesha (nnesha):

6 times 6 ? there is only one 6

Nnesha (nnesha):

The variable Z is directly proportional to X z= kx and 2nd part is inversely proportional to Y so z =kx/y

Nnesha (nnesha):

substitute x,y and z values solve for k

OpenStudy (anonymous):

so what do i do first?

Nnesha (nnesha):

substitute x ,y and z values solve for k

OpenStudy (anonymous):

im so confused

Nnesha (nnesha):

why ? what you didn't understand :=)

OpenStudy (anonymous):

i substitued and got 10 and 16

Nnesha (nnesha):

so how did you get 10 please show your work so i can figure t out ur mistake :=)

OpenStudy (anonymous):

oh wait no , i got 60 and 24

Nnesha (nnesha):

work ?

Nnesha (nnesha):

k= ?

OpenStudy (anonymous):

all i did was multiply 10 by 6 and 8 by 3

Nnesha (nnesha):

When X is 6 and Y is 3, Z has the value 6. this is first part you need to find k value first ALWAYS!

Nnesha (nnesha):

substitute x for 6 y for 3 z for 6 solve for k